# Selfless reduced free product $C^{*}$-algebras

### v1

@@ 1.1. Context and motivation
Studying strict comparison in $C^{*}$-algebras has by now several demonstrably important applications. To illustrate this, let us fix, for convenience, a simple monotracial $C^{*}$-algebra $A$ with strict comparison. We can then explicitly calculate the Cuntz semigroup of $A$ (e.g., see [3]). We know also, by a recent theorem of Lin, that $A$ must have stable rank one [35]. Further, by the work [46], we can fully classify the embeddings of the Jiang-Su algebra $\mathcal{Z}$ into $A$ up to approximate unitary equivalence (note also, the upcoming work [11] which deals with a vast generalization). More recently, strict comparison led in [51] to solve the $C^{*}$-algebraic Tarski problem, strikingly in the negative. In particular, it was proved there that the $K_{1}$-group of an ultraproduct of $C^{*}$-algebras can be computed under a strict comparison hypothesis.
@@
In the recent work [47], the third-named author introduced the notion of a selfless $C^{*}$-probability space. Selflessness yields several structural properties in a unified manner: simplicity, at most one tracial state, in the case of a faithful state the dichotomy ‘‘purely infinite or stable rank one’’, and, significantly, strict comparison. An important breakthrough came then in [1], where it was proved that large classes of reduced group $C^{*}$-algebras are selfless, including those arising from acylindrically hyperbolic groups with the rapid decay property and trivial finite radical. This settled affirmatively the question of strict comparison for $C_{\lambda}^{*}(\mathbb{F}_{2})$. The strategy developed in [1] (notably inspired by the key work [36]) has led to a more recent advancement that obtains selflessness and, therefore, strict comparison and stable rank one, for first examples of reduced group $C^{*}$-algebras of higher-rank lattices [57] (see also [5]).
@@ 1.2. Main results
  • $C(G)$ where $G$ be a connected compact Lie group and letting $\tau$ be the state on $C(G)$ induced by the Haar measure.
@@
The conditions on the unitaries above are motivated by, and closely related to, work of Avitzour [4]. Note that a finite-dimensional $C^{*}$-probability space $(A,\rho)$ with faithful state $\rho$ always has the rapid decay property relative to the filtration $V_{n}=A$ for all $n$. We can thus deduce from Theorem B that $(\mathbb{M}_{m}(\mathbb{C}),\rho_{1})*(\mathbb{M}_{n}(\mathbb{C}),\mathrm{tr})$ is selfless once $m,n\geq 2$ and once $\rho_{1}$ is faithful. If $\rho_{1}$ is additionally nontracial, then $(\mathbb{M}_{m}(\mathbb{C}),\rho_{1})*(\mathbb{M}_{n}(\mathbb{C}),\mathrm{tr})$ is purely infinite. Interestingly, just as strict comparison in $C_{\lambda}^{*}(\mathbb{F}_{2})$ had resisted earlier approaches, pure infiniteness of such reduced free products had remained out of reach until the methods of [1], building on [47]. Other examples of selfless reduced free products can be obtained from Theorem B, e.g., $(A,\tau)*({\mathbb{M}}_{n}({\mathbb{C}}),\operatorname{tr})$ if $n\geq 2$, $A$ is finite-dimensional, and $\tau$ is a faithful trace with $\tau(z)<1/2$ for every central projection in $A$.
@@
The second bullet point holds whenever the GNS completion of $A_{2}$ with respect to $\tau$ has diffuse center. Thus, we may apply Theorem C when $A_{2}=C(X)$ and $\tau=\int\cdot\,d\mu$ where $\mu$ is atomless, provided that we have filtrations of $A_{1},A_{2}$ with the rapid decay property. This allows us to deduce that the $C^{*}$-algebra of $n$ free semicirculars is selfless (Theorem A) once $n\geq 2$. As explained before, this case is not directly approachable via Theorem B, since our filtrations for the semicircular measure do not have nonconstant unitaries. Since we also do not require that $A_{1}$ has unitaries with state zero, we can use Theorem C to show that $({\mathbb{C}}^{2},\rho)*(C^{*}_{\lambda}(G),\tau)$ is selfless, when $\rho$ is any faithful state, $G$ is any icc group with the rapid decay property, and $\tau$ is the usual trace from the left regular representation. (This also does not follow from Theorem B.) Note however that Theorem B does imply that free products such as $({\mathbb{M}}_{k}({\mathbb{C}}),\operatorname{tr})*({\mathbb{M}}_{n}({\mathbb{C}}),\operatorname{tr})$ (with $n,k\geq 2$) and ${\mathbb{M}}_{k}({\mathbb{C}})*C^{*}(G)$ (with $k\geq 2$ and $G$ a finite nontrivial group) are selfless. These cases do not directly follow Theorem C, as the assumptions of that theorem imply that one of the sides of the free product has diffuse GNS completion. The two theorems are thus complementary in some sense.
@@
Using both cases of Theorem C together with Dykema’s classification theorems [17, 16], we completely classify selflessness among free products of finite dimensional abelian $C^{*}$-algebras:
@@ 1.3. Roadmap of the proofs
We now discuss Theorem C. The method of proof here again involves maps $\phi_{n}$ of the form $\phi_{v_{n}}$, arising from unitaries $v_{n}$, as above. In the process of adapting the proof strategy for Theorem B to more general free products, we no longer wish to choose these unitaries inside the filtrations of $A_{1}$ and $A_{2}$. Instead, we will choose $v_{n}\in\mathcal{U}(A_{2})$. In order to get the $L^{2}$-$L^{2}$ isometry argument to go through, we choose $(v_{n})_{n}$ satisfying suitable asymptotic orthogonality conditions. This is accomplished in Lemma 6.3. Note that the images under $\phi_{n}$ of the subspaces from the filtration of $B_{1}*_{\textnormal{alg}}B_{2}*_{\textnormal{alg}}B_{3}$ are no longer contained in subspaces from the filtration of $A_{1}*A_{2}$. Thus, the new arguments need to ensure that these images are contained is subspaces of $A_{1}*A_{2}$ where the operator norm is still controlled by the $L^{2}$-norm, i.e., a form of ‘‘inflated rapid decay’’. This step is more subtle. We discuss next in more detail the challenges that it entails.
@@
and the right-hand side above is approximately $C"\|x_{1}v+v^{*}x_{2}+x_{3}\|_{2}^{2}$.
@@ 2.1. Preliminary notation
Recall that if $r\in{\mathbb{N}}$ and $\Omega$ is a set, then a function $j\colon[r]\to\Omega$ is alternating if $j(i)\neq j(i+1)$ for all $11\leq i\leq r-$ 1. If $(A,\rho)$ is a $C^{*}$-probability space, and $S_{1},\cdots,S_{m}\subseteq A$ we say that $a=a_{1}\cdots a_{\ell}$ is an alternating word in $(S_{1},\cdots,S_{m})$ if: $a_{i}\in S_{j(i)}$ for an alternating function $j\colon[\ell]\to[m]$. If, in addition, $\rho(a_{i})=0$ for all $i$, then we call $a_{1}\cdots a_{\ell}$ an alternating centered word in $(S_{1},\cdots,S_{m})$. In our analysis of alternating centered words, we will occasionally have to deal with words of length zero. Note that, by the description of the free product Hilbert space, if $x\in(A_{1},\rho_{1})*\cdots*(A_{m},\rho_{m})$ is an alternating centered word in $(A_{1},\cdots,A_{m})$, then the decomposition of $x$ as an alternating centered word in $(A_{1},\cdots,A_{m})$ is unique. Sometimes, especially in Section 6 , we will deal with alternating words that are not centered. In this case, the decomposition as an alternating word is not unique. Care must be taken in such cases, but this will not present an issue in our proofs. As we will usually be in the unital case, we will take the convention that the empty product in a unital algebra is equal to $1$. Similarly, we take the convention that the empty sum is zero.
@@ 3.1. The rapid decay property
for all $a\in V_{n}$. When $(V_{n})_{n}$ is the degree‐filtration of $C^{*}(x,1)$ coming from a tuple $x\in A^{r}$, we say that the tuple $x$ has rapid decay.
@@
We give $S^{\infty}((V_{n})_{n})$ and $S^{2}((V_{n})_{n})$ a locally convex structure by equipping them with the family of norms $(\|\cdot\|_{S^{\infty,\alpha}})_{\alpha>0}$ and $(\|\cdot\|_{S^{2,\alpha}})_{\alpha>0}$, respectively. Restricting these families of seminorms to $\alpha\in{\mathbb{N}}$ defines the same topology and thus these locally convex topologies are metrizable. These topologies are also complete (the verification is straightforward and is left to the reader). In [60, Remark 5.10, Proposition 5.11], Voiculescu gives multiple characterizations of $S^{\infty}((V_{n})_{n})$.
@@
To see that (v) implies (i), note that $S^{2}((V_{n})_{n})$ is a locally convex Frechet space. By a version of the closed graph theorem valid for Frechet spaces (see [49, Theorem 2.15], we see that the inclusion map $\iota\colon S^{2}((V_{n})_{n})\to A$ is necessarily continuous. By the continuity criteria for linear maps between locally convex spaces ([20, Proposition 5.15]), we see that there is an $\alpha>0$ such that
@@
    Suppose that $A_{2}=C_{\lambda}^{*}(G)$, where $G$ is a finitely generated discrete group with rapid decay, $(V_{n,2})_{n}$ is a filtration on $C_{\lambda}^{*}(G)$ obtained from a finite generating set (see Example 1), and $\rho_{2}$ the canonical trace on $C_{\lambda}^{*}(G)$. Then
@@
(iii) Below $\odot$ denotes the algebraic tensor product and $\otimes$ the minimal one.
Let us assume that $A_{2}=C_{\lambda}^{*}(G)$ and that $\rho_{2}$ is the canonical trace. Let us first prove a claim from which (iii) will readily follows. Let $V\subseteq A_{1}$ and $W\subseteq A_{2}$ be linear subspaces, with $W$ spanned by a finite set of canonical unitaries $\{u_{g}:g\in F\}$. Let us show that if
$\|a\|\leq K\|a\|_{2},\,\|b\|\leq L\|b\|_{2},\quad\forall a\in V,\,\forall b\in W,$
then $\|c\|\leq KL\|c\|_{2}$ for all $c\in V\odot W$. Since $A_{1}$ and $A_{2}$ are faithfully represented in $B(L^{2}(A_{1},\rho_{1}))$ and $B(L^{2}(A_{2},\rho_{2}))$ (by Proposition 3.3), and since we work with the minimal tensor product, this is equivalent to
$\|cd\|_{2}\leq KL\|c\|_{2}\|d\|_{2},\quad\forall c\in V\odot W,\quad\forall d\in A_{1}\odot A_{2}.$
The proof of this inequality is a minor adaptation of [31, Lemma 1].
Given $c,d\in A_{1}\odot A_{2}$, let us write them as
$c=\sum_{g\in G}c_{g}\otimes u_{g},\quad d=\sum_{g\in G}d_{g}\otimes u_{g}$
where $c_{g},d_{g}\in A_{1}$ and $g\in G$. Define $\theta_{c},\theta_{d}\in A_{2}$ by
$\theta_{c}=\sum_{g}\|c_{g}\|_{2}u_{g},\,\theta_{d}=\sum_{g}\|d_{g}\|_{2}u_{g}.$
Notice that $\|\theta_{c}\|_{2}=\|c\|_{2}$, $\|\theta_{d}\|_{2}=\|d\|_{2}$, and that
$\|\theta_{c}\theta_{d}\|_{2}^{2}=\sum_{g}\Big{(}\sum_{hh^{\prime}=g}\|c_{h}\|_{2}\|d_{h^{\prime}}\|_{2}\Big{)}^{2}.$
Assume now that $c\in V\odot W$. Then,
$cd=\sum_{g,g^{\prime}}c_{g}d_{g^{\prime}}\otimes u_{gg^{\prime}}=\sum_{g}\Big{(}\sum_{hh^{\prime}=g}c_{h}d_{h^{\prime}}\Big{)}\otimes u_{g}.$
Here $h$ ranges in $F$ and $c_{h}\in V$ for all $h\in F$. Thus,
$\|cd\|_{2}^{2}$ $=\sum_{g}\Big{\|}\sum_{hh^{\prime}=g}c_{h}d_{h^{\prime}}\Big{\|}_{2}^{2}$
$\leq\sum_{g}\Big{(}\sum_{hh^{\prime}=g}\|c_{h}d_{h^{\prime}}\|_{2}\Big{)}^{2}$
$\leq K^{2}\sum_{g}\Big{(}\sum_{hh^{\prime}=g}\|c_{h}\|_{2}\|d_{h^{\prime}}\|_{2}\Big{)}^{2}$
$=K^{2}\|\theta_{c}\theta_{d}\|_{2}^{2}\leq K^{2}L^{2}\|\theta_{c}\|_{2}^{2}\theta_{d}\|_{2}^{2}=K^{2}L^{2}\|c\|_{2}^{2}\|d\|^{2}_{2}.$
We thus deduce that $\|cd\|_{2}\leq KL\|c\|_{2}\|d_{2}\|$.
We now readily verify that if $A_{1}$ and $A_{2}=C_{\lambda}^{*}(G)$ have rapid decay relative to filtrations $(V_{n,1})_{n}$ and $(V_{n,2})_{n}$, respectively (the latter as defined in Example 1), and $\|a\|\lesssim(1+n)^{\alpha}\|a\|_{2}$ for $a\in V_{n,1}$ and $\|b\|\lesssim(1+n)^{\beta}\|b\|_{2}$ for $b\in V_{n,2}$, then
$\|c\|\lesssim(1+n)^{\alpha+\beta}\|c\|_{2},\quad\forall c\in V_{n,1}\odot V_{n,2}.$

### v2

@@ 1.1. Context and motivation
Studying strict comparison in $C^{*}$-algebras has by now several demonstrably important applications. To illustrate this, let us fix, for convenience, a simple monotracial $C^{*}$-algebra $A$ with strict comparison. We can then explicitly calculate the Cuntz semigroup of $A$ (e.g., see [3]). We know also, by a recent theorem of Lin, that $A$ must have stable rank one [38]. Further, by the work [49], we can fully classify the embeddings of the Jiang-Su algebra $\mathcal{Z}$ into $A$ up to approximate unitary equivalence (note also, the upcoming work [11] for related results). More recently, strict comparison led in [54] to solve the $C^{*}$-algebraic Tarski problem, strikingly in the negative. In particular, it was proved there that the $K_{1}$-group of an ultraproduct of $C^{*}$-algebras can be computed under a strict comparison hypothesis.
@@
In the recent work [50], the third-named author introduced the notion of a selfless $C^{*}$-probability space. Selflessness yields several structural properties in a unified manner: simplicity, at most one tracial state, if the state is faithful the dichotomy ‘‘purely infinite or stable rank one’’, and, significantly, strict comparison. An important breakthrough came then in [1], where it was proved that large classes of reduced group $C^{*}$-algebras are selfless, including those arising from acylindrically hyperbolic groups with the rapid decay property and trivial finite radical. This settled affirmatively the question of strict comparison for $C_{\lambda}^{*}(\mathbb{F}_{2})$. The strategy developed in [1] (notably inspired by the key work [39]) has led to a more recent advancement that obtains selflessness and, therefore, strict comparison and stable rank one, for first examples of reduced group $C^{*}$-algebras of higher-rank lattices [62] (see also [5]).
@@ 1.2. Main results
  • $C(G)$ where $G$ is a connected compact Lie group and letting $\tau$ be the state on $C(G)$ induced by the Haar measure.
@@
The conditions on the unitaries above are motivated by, and closely related to, work of Avitzour [4]. Note that a finite-dimensional $C^{*}$-probability space $(A,\rho)$ with faithful state $\rho$ always has the rapid decay property relative to the filtration $V_{n}=A$ for all $n$. We thus deduce from Theorem B that
$(\mathbb{M}_{m}(\mathbb{C}),\rho_{1})*(\mathbb{M}_{n}(\mathbb{C}),\mathrm{tr})$
is selfless whenever $m,n\geq 2$ and $\rho_{1}$ is faithful. If $\rho_{1}$ is additionally nontracial, then this reduced free product is purely infinite. We note that Theorem B (and Theorem C below), provide relatively simple criteria for proving that a reduced free product is purely infinite (by establishing selflesness relative to a faithful nontracial state). Previous results (for example, [19]) established pure infiniteness of reduced free products in certain cases, under more complicated hypotheses.
Other examples of selfless reduced free products can be obtained from Theorem B, e.g., $(A,\tau)*({\mathbb{M}}_{n}({\mathbb{C}}),\operatorname{tr})$ if $n\geq 2$, $A$ is finite-dimensional, and $\tau$ is a faithful trace with $\tau(z)<1/2$ for every central projection in $A$.
@@
The second bullet point holds whenever the GNS completion of $A_{2}$ with respect to $\tau$ has diffuse center. Thus, we may apply Theorem C when $A_{2}=C(X)$ and $\tau=\int\cdot\,d\mu$ where $\mu$ is atomless, provided that we have filtrations of $A_{1},A_{2}$ with the rapid decay property. This allows us to deduce that the $C^{*}$-algebra of $n$ free semicirculars is selfless (Theorem A) once $n\geq 2$. As explained before, this case is not directly approachable via Theorem B, since our filtrations for the semicircular measure do not have nonconstant unitaries. We then deduce from Theorem A, that the free Araki–Woods C*-algebras $\Gamma(H_{\mathbb{R}},U_{t})$ introduced by Shylakhtenko are selfless, simple, and purely infinite, for $\dim H_{\mathbb{R}}\geq 3$.
@@
Since we do not require that $A_{1}$ has unitaries with state zero, we can use Theorem C to show that
$({\mathbb{C}}^{2},\rho)*(C^{*}_{\lambda}(G),\tau)$
is selfless, when $\rho$ is any faithful state, $G$ is any icc group with the rapid decay property, and $\tau$ is the usual trace from the left regular representation. This does not follow from Theorem B. Note however that Theorem B does imply that free products such as
$({\mathbb{M}}_{k}({\mathbb{C}}),\mathrm{tr})*(C^{*}_{\lambda}(G),\tau),$
with $k\geq 2$ and $G$ a finite nontrivial group, are selfless. These cases do not directly follow Theorem C, as the assumptions of that theorem imply that one of the sides of the free product has diffuse GNS completion. The two theorems are thus complementary in some sense.
Using both cases of Theorem C, we completely classify selflessness among free products of finite dimensional abelian $C^{*}$-algebras:
@@
Besides Theorem C, to prove Theorem D we rely on Dykema’s structure result on the reduced free products of finite-dimensional abelian C*-algebras from [18], as well as on Dykema’s classification of von Neumann free products of finite-dimensional abelian von Neumann algebras [16]. We mention that if $(M_{1},\rho_{1})$ and $(M_{2},\rho_{2})$ are von Neumann algebras equipped with faithful, normal, states, then [17, Theorem 3] determines precisely when their von Neumann free product is a type III factor (see [58, Theorem 3] for an alternative approach). Furthermore, a rather surprising result of Hartglass-Nelson [30] precisely classifies the free products of finite-dimensional von Neumann algebras in terms of the Free Araki–Woods factors of Shlyakhtenko [55]. It is plausible that these results can be used to investigate selflessness of the reduced free product of two finite-dimensional $C^{*}$-probability spaces. However, adapting our approach to this setting poses substantial challenges. We leave this question as an interesting problem for future research.
@@ 1.3. Roadmap of the proofs
We now discuss Theorem C. The method of proof here again involves maps $\phi_{n}$ of the form $\phi_{v_{n}}$, arising from unitaries $v_{n}$, as above. In the process of adapting the proof strategy for Theorem B to more general free products, we no longer wish to choose these unitaries inside the filtrations of $A_{1}$ and $A_{2}$. Instead, we will choose $v_{n}\in\mathcal{U}(A_{2})$. In order to get the $L^{2}$-$L^{2}$ isometry argument to go through, we choose $(v_{n})_{n}$ satisfying suitable asymptotic orthogonality conditions. This is accomplished in Lemma 6.3. Note that the images under $\phi_{n}$ of the subspaces from the filtration of $B_{1}*_{\textnormal{alg}}B_{2}*_{\textnormal{alg}}B_{3}$ are no longer contained in subspaces from the filtration of $A_{1}*A_{2}$. Thus, the new arguments need to ensure that these images are contained in subspaces of $A_{1}*A_{2}$ where the operator norm is still controlled by the $L^{2}$-norm, i.e., a form of ‘‘inflated rapid decay’’. This step is more subtle. We discuss next in more detail the challenges that it entails.
@@
and the right-hand side above is approximately $C^{\prime\prime}\|x_{1}v+v^{*}x_{2}+x_{3}\|_{2}^{2}$.
@@ 2.1. Preliminary notation
Recall that if $r\in{\mathbb{N}}$ and $\Omega$ is a set, then a function $j\colon[r]\to\Omega$ is alternating if $j(i)\neq j(i+1)$ for all $11\leq i\leq r-$ 1. If $(A,\rho)$ is a $C^{*}$-probability space, and $S_{1},\cdots,S_{m}\subseteq A$ we say that $a=a_{1}\cdots a_{\ell}$ is an alternating word in $(S_{1},\cdots,S_{m})$ if: $a_{i}\in S_{j(i)}$ for an alternating function $j\colon[\ell]\to[m]$. If, in addition, $\rho(a_{i})=0$ for all $i$, then we call $a_{1}\cdots a_{\ell}$ an alternating centered word in $(S_{1},\cdots,S_{m})$. Note that, by the description of the free product Hilbert space, if $x\in(A_{1},\rho_{1})*\cdots*(A_{m},\rho_{m})$ is an alternating centered word in $(A_{1},\cdots,A_{m})$, then the decomposition of $x$ as an alternating centered word in $(A_{1},\cdots,A_{m})$ is unique modulo scaling. Sometimes, especially in Section 6 , we will deal with alternating words that are not centered. In this case, the decomposition as an alternating word is not unique. Care must be taken in such cases, but this will not present an issue in our proofs. In our analysis of alternating centered words, we will occasionally have to deal with words of length zero. As we will usually be in the unital case, we will take the convention that the empty product in a unital algebra is equal to $1$. Similarly, we take the convention that the empty sum is zero.
@@ 3.1. The rapid decay property
for all $a\in V_{n}$. When $(V_{n})_{n}$ is the degree-filtration of $C^{*}(x,1)$ coming from a tuple $x\in A^{r}$, we say that the tuple $x$ has rapid decay.
@@
We give $S^{\infty}((V_{n})_{n})$ and $S^{2}((V_{n})_{n})$ a locally convex structure by equipping them with the family of norms $(\|\cdot\|_{S^{\infty,\alpha}})_{\alpha>0}$ and $(\|\cdot\|_{S^{2,\alpha}})_{\alpha>0}$, respectively. Restricting these families of seminorms to $\alpha\in{\mathbb{N}}$ defines the same topology and thus these locally convex topologies are metrizable. These topologies are also complete (the verification is straightforward and is left to the reader). In [65, Remark 5.10, Proposition 5.11], Voiculescu gives multiple characterizations of $S^{\infty}((V_{n})_{n})$.
@@
To see that (v) implies (i), note that $S^{2}((V_{n})_{n})$ is a locally convex Frechet space. By a version of the closed graph theorem valid for Frechet spaces (see [52, Theorem 2.15], we see that the inclusion map $\iota\colon S^{2}((V_{n})_{n})\to A$ is necessarily continuous. By the continuity criteria for linear maps between locally convex spaces ([22, Proposition 5.15]), we see that there is an $\alpha>0$ such that
@@
    Suppose that the following condition holds: for all $n\in{\mathbb{N}}$, $k\in{\mathbb{N}}$, and all $s\in\{n,k,n+k\}$, there are orthonormal bases $(x_{i,s})_{i\in I_{s}}$ for $\overline{V_{s,2}}^{\|\cdot\|_{2}}$ satisfying that $\langle x_{i,n}x_{j,k},x_{p,n+k}\rangle\geq 0$ for all $i\in[n]$, $j\in[k]$, and $p\in[n+k]$. Then
@@
(iii) By Corollary 3.7, we may assume that $V_{n,2}$ is $\|\cdot\|_{2}$-closed for all $n$. Below $\odot$ denotes the algebraic tensor product and $\otimes$ the minimal one. Choose $\alpha_{j}>0$, with $j=1,2$, so that
$\|x\|\lesssim(n+1)^{\alpha_{j}}\|x\|_{2},\textnormal{ for all $x\in V_{n,j}$ and $j=1,2$.}$
Fix $n,k\in{\mathbb{N}}$, and let $(x_{i,s})_{s\in I_{s}}$ be orthonormal bases as in the assumption for $s\in\{n,k,n+k\}$. Set $E_{s,2}=\operatorname{span}\{x_{i,s}:i\in I_{s}\}$. Let $c\in V_{n,1}\odot E_{n,2}$ and $d\in V_{k,1}\odot E_{k,2}$. Write
$c=\sum_{i\in I_{n}}c_{i}\otimes x_{i,n},\quad d=\sum_{j\in I_{k}}d_{j}\otimes x_{j,k}.$
Let $S_{c}=\{i:c_{i}\neq 0\}$, $S_{d}=\{j:d_{j}\neq 0\}$, and note that both these sets are finite. For $p\in[n+k]$, let $\lambda_{i,j,p}=\langle x_{i,n}x_{j,k},x_{p,n+k}\rangle$. Then:
$cd=\sum_{p}\left(\sum_{(i,j)\in S_{c}\times S_{d}}\lambda_{i,j,p}c_{i}d_{j}\right)\otimes x_{p,n+k}.$
Since $S_{c},S_{d}$ are finite sets, there are no convergence issues when we write the sum above, and the sum over $p$ converges in $\|\cdot\|_{2}$. Set
$\Theta_{c}=\sum_{i\in S_{c}}\|c_{i}\|_{2}x_{i,n},\quad\Theta_{d}=\sum_{j\in S_{d}}\|d_{j}\|_{2}x_{j,k}.$
Using that $\lambda_{i,j,p}\geq 0$ for all $i\in[n],j\in[k],p\in[n+k]$, we have:
$\|cd\|_{2}^{2}=\sum_{p}\left\|\sum_{(i,j)\in S_{c}\times S_{d}}\lambda_{i,j,p}c_{i}d_{j}\right\|_{2}^{2}$ $\leq\sum_{p}\left(\sum_{(i,j)\in S_{c}\times S_{d}}\lambda_{i,j,p}\|c_{i}d_{j}\|_{2}\right)^{2}$
$\lesssim(n+1)^{2\alpha_{1}}\sum_{p}\left(\sum_{(i,j)\in S_{c}\times S_{d}}\lambda_{i,j,p}\|c_{i}\|_{2}\|d_{j}\|_{2}\right)^{2}$
$=(n+1)^{2\alpha_{1}}\|\Theta_{c}\Theta_{d}\|_{2}^{2},$
with the last inequality following by the assumed rapid decay property of $A_{1}$ with respect to $(V_{\ell,1})_{\ell}$. By the assumed rapid decay property of $A_{2}$ with respect to $(V_{t,2})_{t}$, we have that
$\|\Theta_{c}\Theta_{d}\|_{2}\lesssim(n+1)^{\alpha_{2}}\|\Theta_{c}\|_{2}\|\Theta_{d}\|_{2}=(n+1)^{\alpha_{1}}\|c\|_{2}\|d\|_{2}.$
Thus
$\|cd\|_{2}\lesssim(n+1)^{\alpha_{1}+\alpha_{2}}\|c\|_{2}\|d\|_{2}$
for all $c\in V_{n,1}\odot E_{n,2},d\in V_{k,1}\odot E_{k,2}$. By continuity, the same estimate holds for $c\in\overline{V_{n,1}\odot E_{n,1}}^{\|\cdot\|}$ and $d\in\overline{V_{k,1}\odot E_{n,2}}^{\|\cdot\|}$. By Lemma 3.4, it follows that the above estimate holds for $c\in V_{n,1}\otimes V_{n,2}$ and $d\in V_{k,1}\otimes V_{k,2}$. We thus conclude by applying Proposition 3.6. ∎
Note that the hypotheses of (iii) hold if $A_{2}=C^{*}_{\lambda}(G)$, where $G$ is a finitely generated group, or more generally a finitely generated discrete hypergroup (in the sense of [34]), $\rho_{2}$ is the canonical trace on $C_{\lambda}^{*}(G)$, and $(V_{n,2})_{n}$ is the filtration obtained from a finite generating set (see Example 1).